# ISWC 2025: The 24th International Semantic Web Conference, November 2-6, 2025, Nara, Japan

### v1

@@
Copyright for this paper by its authors. Use permitted under Creative Commons License Attribution 4.0 International (CC BY 4.0).
@@
ISWC 2025: The 24th International Semantic Web Conference, November 2-6, 2025, Nara, Japan
@@ 1 Introduction
The evolution of manufacturing systems towards Industry 4.0, bringing flexibility of production processes, re-configurability of production system resources, and better use of industrial artificial intelligence, has fundamentally transformed the operational landscape of industrial production environments [1]. Industrial cyber-physical production systems conforming to Industry 4.0 represent software-intensive architectures, where industrial workcells, robots, and other resources incorporate decisive roles in system functionality, encompassing robot control programs, as well as Manufacturing Execution Systems or Manufacturing Operations Management systems that orchestrate production at the system level [2]. This transformation from centralized to distributed system architectures introduces unprecedented levels of flexibility, enabling mass customization and adaptive manufacturing [3, 4]. As a consequence, production engineering has to shift from a component viewpoint to a system viewpoint and to consider holistic principles of system engineering and analysis [5].
However, such an increased flexibility brought by Industry 4.0 or smart manufacturing presents significant challenges to traditional quality assurance and error diagnostic methodologies [6]. Unlike conventional manufacturing systems that relied on repeatable production sequences executed on identical machines with standardized processes, modern cyber-physical production systems (CPPSs) exhibit dynamic behavior patterns that complicate the testing or commissioning processes, fault detection, and quality control on the system level. The inherent variability in flexible manufacturing makes traditional statistical process control methods inadequate for comprehensive quality management as products are frequently produced on different machines/resources with various production processes, making production processes hard to repeat and significantly more difficult to test and fine-tune. As flexible use cases make it no longer efficient or even feasible to rely on traditional Ishikawa/fishbone diagrams [5], industrial practitioners require more versatile, flexible, and dynamic models and approaches for operations support and quality assurance.
To cope with all these challenges and constraints, we designed a semantic model realizing the product–process–resource (PPR) asset knowledge graph (AKG). The proposed combined model fulfills the following requirements that are based on recommendations from industrial practitioners:
@@
The motivation for the PPR-AKG approach comes from lessons learned in one of our projects: automated robotic disassembly of used electric vehicle batteries and their remanufacturing [7] for stationary energy storages. This particular use case involves highly variable input product and material states, affecting subsequent processes, and complex quality assessment procedures.
@@
The proposed approach “Product–Process–Resource Asset Knowledge Graphs” (PPR-AKGs) extends a Product–Process–Resource (PPR) model, established in [8] and grounded in VDI 3682, ISA-95, and IEC 62264. The PPR-AKG provides a foundational manufacturing system model [9] by recognizing triadic relationship between products, production processes, and production resources for comprehensive system understanding. However, the traditional representations lack the identification of undesired and desired conditions, which are crucial for practical operation in industrial manufacturing. The proposed PPR-AKG model shall close this gap by extending the PPR model with undesired conditions and their causes. In addition, traditional PPR implementations lack the semantic richness necessary for expressing complex dependencies and causal relationships that characterize modern CPPS operations [10]. Therefore, the comprehensive model is represented as OWL (i.e., Web Ontology Language) ontology, benefiting from ontology features such as formal representation, querying, and reasoning.
@@
For industrial practice and to meet Req. 2, it is very important to focus on the undesired conditions (purple boxes) as well as their plausible causes (gray boxes). Plausible causes can be expressed resource-specifically (i.e., inside yellow boxes, which can be defined by the resource vendor), as well as on the system level globally (right-hand column in Fig. 1). Such a scoping of causes for undesired conditions has been recommended by practitioners and it is also inspired by procan.do method [11].

### v2

@@ 1 Introduction
The evolution of manufacturing systems towards Industry 4.0, bringing flexibility of production processes, re-configurability of production system resources, and better use of industrial artificial intelligence, has fundamentally transformed the operational landscape of industrial production environments [Naeem2024]. Industrial cyber-physical production systems conforming to Industry 4.0 represent software-intensive architectures, where industrial workcells, robots, and other resources incorporate decisive roles in system functionality, encompassing robot control programs, as well as Manufacturing Execution Systems or Manufacturing Operations Management systems that orchestrate production at the system level [Novak2022PyMES]. This transformation from centralized to distributed system architectures introduces unprecedented levels of flexibility, enabling mass customization and adaptive manufacturing [Lunsch2022, Wally2021TASE]. As a consequence, production engineering has to shift from a component viewpoint to a system viewpoint and to consider holistic principles of system engineering and analysis [Winkler2021Risk].
However, such an increased flexibility brought by Industry 4.0 or smart manufacturing presents significant challenges to traditional quality assurance and error diagnostic methodologies [Wehner2023Interactive]. Unlike conventional manufacturing systems that relied on repeatable production sequences executed on identical machines with standardized processes, modern cyber-physical production systems (CPPSs) exhibit dynamic behavior patterns that complicate the testing or commissioning processes, fault detection, and quality control on the system level. The inherent variability in flexible manufacturing makes traditional statistical process control methods inadequate for comprehensive quality management as products are frequently produced on different machines/resources with various production processes, making production processes hard to repeat and significantly more difficult to test and fine-tune. As flexible use cases make it no longer efficient or even feasible to rely on traditional Ishikawa/fishbone diagrams [Winkler2021Risk], industrial practitioners require more versatile, flexible, and dynamic models and approaches for operations support and quality assurance.
To cope with all these challenges and constraints, we designed a semantic model realizing the product–process–resource (PPR) asset knowledge graph (AKG). The proposed combined model fulfills the following requirements, substantiated by recommendations from industrial partner practitioners:
@@
The PPR-AKG approach has been used for handing our laboratory prototype over to an industrial vendor in a project: Automated robotic disassembly of used electric vehicle batteries and their remanufacturing [Chigbu2024EVBattRemanufacturing] for stationary energy storages [Strakosova2024ETFAPoPAN]. This use case involves highly variable input product and material states, affecting subsequent processes, and complex quality assessment procedures.
@@
The proposed approach “Product–Process–Resource Asset Knowledge Graphs” (PPR-AKGs) extends a Product–Process–Resource (PPR) model, established in [Schleipen2009] and grounded in VDI 3682, ISA-95, and IEC 62264. The PPR-AKG provides a foundational manufacturing system model [Biffl2021] by recognizing triadic relationship between products, production processes, and production resources for comprehensive system understanding. However, the traditional representations lack the identification of undesired and desired conditions, which are crucial for practical operation in industrial manufacturing. The proposed PPR-AKG model shall close this gap by extending the PPR model with undesired conditions and their causes. In addition, traditional PPR implementations lack the semantic richness necessary for expressing complex dependencies and causal relationships that characterize modern CPPS operations [Terkaj2022]. Therefore, the comprehensive model is represented as OWL (i.e., Web Ontology Language) ontology, benefiting from ontology features such as formal representation, querying, and reasoning.
@@
For industrial practice and to meet Req. 2, it is very important to focus on the undesired conditions (purple boxes) as well as their plausible causes (gray boxes). Plausible causes can be expressed resource-specifically (i.e., inside yellow boxes, which can be defined by the resource vendor), as well as on the system level globally (right-hand column in Fig. 1). Such a scoping of causes for undesired conditions has been recommended by practitioners and it is also inspired by procan.do method [biffl2025introduction].